Jesse doubled a number and then added 5 to get 4l. If he had first added 5
to the original number and then doubled the result, the answer would have been
(a) 18 (b 41 c) 46 (d) 49 (e) 92

T

Answer by amnd(23)

Let's call the original number A:
Double the original number (A), and add 5 to it to gain 41:
2A + 5 = 41
2A = 41 - 5
2A = 36
A = 18
.
So now that we know the value of A, we can find the value asked by the problem:
.
Add 5 to A, then double the result:
.
(A + 5)*2 = (18 + 5)*2 = 23*2 = 46 (c)